Lilian9667:

What is the value of m? 12 + 7 + m = 30 Drag and drop the correct answer in the box. m = Response area A. 11 B. 19 C. 25 D. 35 E. 49

stormyfurr:

The answer is A. 11

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @stormyfurr The answer is A. 11 \(\color{#0cbb34}{\text{End of Quote}}\) You sure?

sylvester:

12 + (7 + m) = 30 m = 30 - 12 - 7 = 11


12+7+11= 30

supie:

yea

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie yea \(\color{#0cbb34}{\text{End of Quote}}\) She's right

sylvester:

Answer 11

Lilian9667:

Ok Thanks

xxcoffee:

its 11

Dmoney23:

not trying to be rude but stop giving direct answers

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxcoffee its 11 \(\color{#0cbb34}{\text{End of Quote}}\) Thank you

xxcoffee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 \(\color{#0cbb34}{\text{Originally Posted by}}\) @xxcoffee its 11 \(\color{#0cbb34}{\text{End of Quote}}\) Thank you \(\color{#0cbb34}{\text{End of Quote}}\) np

xxcoffee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Dmoney23 not trying to be rude but stop giving direct answers \(\color{#0cbb34}{\text{End of Quote}}\) she/he asked for help so we gave the right anwser

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxcoffee \(\color{#0cbb34}{\text{Originally Posted by}}\) @Dmoney23 not trying to be rude but stop giving direct answers \(\color{#0cbb34}{\text{End of Quote}}\) she/he asked for help so we gave the right anwser \(\color{#0cbb34}{\text{End of Quote}}\) Yah

supie:

yeah but not everyone gave a direct answer tho

supie:

oh nvm everyone did cep for Sylvester

supie:

accept*

Dmoney23:

yes they did everybody's response was said 11

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Dmoney23 yes they did everybody's response was said 11 \(\color{#0cbb34}{\text{End of Quote}}\) no sylvester didn't

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie @Sylvester oh nvm everyone did cep for Sylvester \(\color{#0cbb34}{\text{End of Quote}}\) Yah Sylvest didn't give Direct answer

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Dmoney23 yes they did everybody's response was said 11 \(\color{#0cbb34}{\text{End of Quote}}\) Cus that's the Answer

sylvester:

yesss siirrr lol

xxcoffee:

but she/he did aask so we gave it

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sylvester 12 + (7 + m) = 30 m = 30 - 12 - 7 = 11 12+7+11= 30 \(\color{#0cbb34}{\text{End of Quote}}\) this isn't direct

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxcoffee but she/he did aask so we gave it \(\color{#0cbb34}{\text{End of Quote}}\) Exackly

Dmoney23:

yes it is if it has 11 in it it is direct

angiesyl:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie oh nvm everyone did cep for Sylvester \(\color{#0cbb34}{\text{End of Quote}}\) lmao true

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Dmoney23 yes it is if it has 11 in it it is direct \(\color{#0cbb34}{\text{End of Quote}}\) Boi you confusing

sylvester:

ugh yall makin a big deal for wha

xxcoffee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Dmoney23 yes it is if it has 11 in it it is direct \(\color{#0cbb34}{\text{End of Quote}}\) Boi you confusing \(\color{#0cbb34}{\text{End of Quote}}\) he is like omg

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sylvester ugh yall makin a big deal for wha \(\color{#0cbb34}{\text{End of Quote}}\) yeah frfr

Dmoney23:

yall keep doing what yall are doing, not my fault if you get a warning or suspended

Lilian9667:

ALLRIGHT GUYS

Lilian9667:

I'MA CLOSE IT OK

supie:

ight

Lilian9667:

I GOT THE ANSWER IS DONE

supie:

ighty